In the following first order completing reaction.
A + Reagent ----Product
B + Reagent-----Product
The ratio of K1/K2 if only 50% of B will have been reacted,When 94% of A has been reacted is
 

Grade:12

1 Answers

Umakant biswal
7 years ago
@ ebin 
i can help u with the correct approach to this question , rest u have to only put value and solve it . 
apply the following formula 
K = 2.303 / t log [ initial concentration } / [ final concentration ] 
in the same way u need to calculate k1 and k2 
in the frst equation 
k1 = 2.303 / t log 100 / 50 
k2 = 2.303 / t log 100 / 6 
and then take the ratio 
”HOPE IT CLEARS NOW